How many milliliters of a 7% alcohol solution must be added to 5 milliliters of a 31% alcholol solution to produce a 15% alcohol solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 31 ---------------- 5 ml
Alcohol II 7 ---------------- x ml
Mixture 15 ---------------- 5 + x ml

31*5+7x =15(5+x )
155+7x = 75 + 15 x
7x-15x = 75 - -155
-8x= -80
/ -8
x = 10
10 ml of 7 % Alcohol II